OpenStudy (anonymous):

6(g + 3) = -2(g + 31) a. -2 c. 2.9 b. -10 d. 6

OpenStudy (ivettef365):

first solve the parenthesis

OpenStudy (shamim):

\[6g+18=-2g-62\]

OpenStudy (shamim):

\[8g=-80\]

OpenStudy (goldphenoix):

Do you know how to do the distributive property?

OpenStudy (ivancsc1996):

The divide by 8

OpenStudy (shamim):

ya

OpenStudy (anonymous):

i get all tht but i still dont get the anwser

OpenStudy (goldphenoix):

6g + 1

OpenStudy (goldphenoix):

Does it make sense so far?

OpenStudy (anonymous):

mhm

OpenStudy (goldphenoix):

Alright. So your equation so far looks like: 6g + 18 = -2g - 62 What do you think do next?

OpenStudy (anonymous):

do the two number with the like terms g? 6g and -2g?

OpenStudy (goldphenoix):

Okay. So after doing that, what does your equation look like?

OpenStudy (anonymous):

either 4g+18= -62 or 8g+18 =-62 i think 0.0

OpenStudy (goldphenoix):

So you want to cancel out -2, right? So how can we cancel out -2?

OpenStudy (anonymous):

blehh im lost

OpenStudy (goldphenoix):

6g + 18 = -2g - 62 What do we do? How can we find g?

OpenStudy (anonymous):

nvm im not gunna do this one nomore

OpenStudy (goldphenoix):

o.O? I want you to understand it.

OpenStudy (goldphenoix):

I'm going to explain it to you like a teacher.

OpenStudy (goldphenoix):

So after doing the distributive property, your equation should look like: 6g + 18 = -2g - 62 We want to find g. Though we have 2 g in this equation. One on each side. We want one g, to find g. (If that make sense) So we must get rid of 1 g. I can subtract 6g or add 2g. I'm going to add 2g because it's easier if the number is positive, not negative. So after adding 2g to both side, your equation should look like: 8g + 18 = - 62 We want to balance the number with variable and the number without variable. If we add 62 to both side, then we can never find what g is. So we must subtract 18 to both side. After subtracting, your equation should look like: 8g = -80. Our final step to find g, is to divide 8. 8g divided by 8 is equal to g. (Cancel out) -80 divided by 8 = -10. So the answer is g = -10. Make sense? I hope this help! If you want to make sure that your answer is correct, you can plug it in. 6(g + 3) = -2(g + 31) would become: 6(-10 + 3) = -2(-10 + 31) Do the parenthesis, -10 + 3 = -7. -10 + 31 = -21. So your equation should look like: 6(-7) = -2(-21) Multiply: 6 * -7 = -42 -2 * -21 = -42 -42 = -42 This mean that g is -10. :)

OpenStudy (anonymous):

ohhhhhh u coulda been said tht i understand lmaoo